24.1 KJ is used to heat a piece of aluminium of mass 250 g. What will the final temperature be if its initial temperature is 15.1°C? The specific heat of aluminium is 0.902 J/g.K.
1
Expert's answer
2020-05-19T08:57:56-0400

Let final temperature be "T" ,

Net heat supplied"=mc\triangle T"

"24.1\ KJ=24100\ J=250\times 0.902\times (T-15.1)"

"T-15.1=106.87\degree C\implies T=121.87\degree C"